Biosphere 2 Research Wiki » xe.blog.code.title » Blog categories » Data-logging and Controls

Data-logging and Controls

Last modified by Michael Sibayan on 2015/02/13 12:32
Comments (0) · Attachments (0) · History

Category: Data-logging and Controls (28 posts) [RSS]

Sep 16 2019

1 sec/min Data logging Performance Issue

There was a performance issue with the main dsc.b2 server with generating log files for 1 minute and 1 second log files.

Affected time range: 9/11/2019 8:10 PM to 9/16/2019 10:40AM

Oct 30 2018

PE102-038 East Bay -2 Swap

The PE102-038 in East bay -2 position was measuring about 1/2 of expected counts. Flowmeter from the +4 position in East bay was swapped with -2 position until a replacement is available. 

Sep 26 2018

Added UV sensor to LEO-OWS

A SU-100-SS-L-20 was added to the LEO-OWS.

*Found that 15 minute moving average for the HMP60, SQ110, LP02, was incorrect and has been fixed.
Unsure if 15 minute moving averaged data before 9/26/2018 12:15pm is correct.

Sep 11 2017

Tipping buckets LEO

Updated- 3/6/2019

West -1 NovaLynx tipping bucket was broken, Shell was replaced on 9/8/17. Calibration done and changed in DSM on server 9/11/17 12:45 to .033305 L/tip

East-2 NovaLynx tipping bucket was re-wired, possible bad connection creating spikes in data, will monitor over next rains and re-evaluate

1/15/19- Replaced reed switches on East Bay Nova Lynx +1 and +2 locations due to switch bouncing and noisy data. 


**The Nova Lynx proxy sensors have been rewired to eliminate the terminal block inside the tipping bucket enclosures. Rusting of the terminal blocks' terminals have been shorting the terminals causing noise. The voltage suppressor bi-directional diodes have been moved out of the rusting terminal blocks to the terminal blocks in the enclosure. Note that the terminal blocks were left inside the enclosures.

2/19/2019 - Rewired LEO East
3/6/2019 - Rewired LEO West. 3/13/2019 - LEO West leak tipping bucket

3/7/2019 - Rewired LEO Center bay

Dec 06 2016

2016126100540 East +4 flow meter swap

On Dec. 5, 2016 from 13:45-13:55, Nate changed out the PE102-75 flow meter with a PE102-38. Looking at the raw counter data it seems to be working. This was a used ebay purchase. 

Data before this time used incorrect scaling factor (1000, instead of the correct 500 for a -75).

Scaling factor was not changed at this point, since the correct value is already in place.

Dec 05 2016

2016125113825 Tipping buckets using incorrect scaling factors

We discovered today the drainage tipping buckets on the East and Center slopes are using the wrong scaling factor for output data. This only effects scaled data.

Nov 21 2016

Instrumentation panel damage - West P03

At approximately 9:30 on Nov. 21 an instrumentation panel (West-P03) was damaged along with a few sensor cables. Repairs underway.

Nov 15 2016

PE102-075 flow meter added to Center Bay Leak collector

Added a Seametrics PE102-075 flow meter inline with the Leak tipping bucket in the Center Bay. 

Valid flow rates for this model are 0.2 - 20 GPM. 

Data for Leak tipping bucket between 11/15/2016 11:40 - 13:06 are not valid

Nov 14 2016

Seepage clogged

Seepage clogged and fixed,flow rates down while fixing ...

Nov 07 2016

Irrigation valve actuator swap 2016117093157

On Nov. 6, 2016 from about 10:45 PM - 11:30 PM Aaron and Michael performed work on the electric actuators controlling/monitoring LEO irrigation. 

Several of the actuators were not responding (No diagnostics have been performed), so we swapped around actuators such that each bay now should have a functional actuator on both the MULTI1 and MULTI2 circuit. 

Due to this change several irrigation circuits will display as "in transition" because they have been physically removed / electrically disconnected.

Tags:
Created by Michael Sibayan on 2015/02/13 12:32

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation